# Muffled Murmurs: Environmental effects in the LISA stochastic signal from stellar-mass black hole binaries

### v1

@@ IV.3 Measurability of dynamical friction effects from a sub-population of sBBHs
Figure 8: SGWB spectra $\Omega_{\mathrm{GW}}(f)$ for various environmental fractions $p_{\mathrm{env}}$ in the case of dynamical friction with $\rho^{\rm inj}=10^{-10}\mathrm{g\,cm^{-3}}$. The gray dashed and black dotted curves indicate the power-law sensitivity (PLS) and Bayesian power-law sensitivity (BPLS) of LISA, under the same assumptions in Figs. 2 and 1.
The impact of the SGWB due to environmental effects will depend on the distribution of the environmental parameters ($\rho$ and $f_{\rm Edd}$) across the population. In a realistic scenario, only a sub-population of stBBH formed in gaseous medium, where the environmental effects are significant. As a proof-of-concept, we phenomenologically construct a mixture model to account for environmental effects arising from a sub-population. We introduce the fraction parameter $p_{\mathrm{env}}$, which characterizes the fractional contribution of the SGWB arising from binaries affected by environmental effects. The total SGWB is then given by
@@
In Fig. 8, for a given injected density of $\rho^{\rm inj}=10^{-10}\mathrm{g}\,\mathrm{cm}^{-3}$, we illustrate how changing the injected fraction $p_{\rm env}^{\rm inj}$ leads to changes in the SGWB. The case of $p_{\rm env}=0$ is identical to the vacuum SGWB, while $p_{\rm env}=1$ is identical to having all binaries affected by environmental effects (as shown in Fig. 1). Notably, for $10<p_{\rm env}<$ 1 the total SGWB exhibits a spectral shape resulting from the $p_{\rm env}$-weighted superposition of two SGWBs, dominated by vacum component at low frequencies. If future LVK observations provide tighter constraints on the binary population, the difference in amplitude of the SGWB at high and low frequencies might signal the presence of a sub-population containing environmental effects.
@@
With only a sub-population affected by environmental effects, we assess how the measurability of $\rho$ changes when varying $p_{\rm env}$. We generate different injections of $\Omega_{\rm Frac}$ using Eq. 29 for $p^{\rm inj}_{\rm env}\in\{0.1,0.3,0.5,0.7,0.9\}$ and a fixed density $\rho^{\rm inj}=10^{-10}\mathrm{g}\,\mathrm{cm}^{-3}$. To perform parameter estimation with $\Omega_{\rm Frac}$, we use the RPLP models given by Eq. 18 for both $\Omega_{\rm vac}$ (with $\alpha=0$) and $\Omega_{\rm env}$. Since we quantified the role of the Galactic foreground in Sec. IV.2.1, we assume its parameters known in this analysis.
We show the one-dimensional marginalized posteriors (orange violins) of $\rho$ in Fig. 9 for the different $p_{\rm env}^{\rm inj}$ considered. For comparison, we also show the one-dimensional marginalized posterior of $\rho$ from Fig. 3, where the injected data contains dynamical friction effects for all binaries with $p_{\rm env}=1$. As the fraction $p_{\rm env}^{\rm inj}$ increases from $0.1$ to $0.9$, the posterior gets better constrained. For each injection, we separately infer with the vacuum model and show the Bayes factors in favor of the mixture model in the top x-axis of Fig. 9. Even with $p_{\rm env}^{\rm inj}=0.1$, the non-vacuum model is confidently preferred. Thus, as a proof-of-concept, we see that LISA is sensitive to the changes in the SGWB caused by a sub-population of sBBHs undergoing dynamical friction.
@@
Figure 9: Marginalized posterior distribution of $\log_{10}(\rho/\rho_{0})$ as a function of the environmental fraction parameter $p_{\rm env}$, with an injected density $\rho^{\rm inj}=10^{-10}\rm{g\,cm^{-3}}$ (horizontal dashed line). The upper x-axis shows the log Bayes factor between the mixture and vacuum models. The gray violin plot, shown for reference, is the posterior recovered using a pure dynamical friction model for an injection with $p_{\rm env}=1$.
@@ V Agnostic tests of additional energy loss channels
In general, when there is energy loss from binaries in addition to GW emission, the SGWB will decrease below the vacuum GR prediction given by Eq. 3. We use our phenomenological model to probe such additional energy losses agnostically. Specifically, with our RPL model (see Eq. 17), one can constrain the dimensional combination $\tilde{\alpha}\equiv A_{m}\alpha$ for different choices of $\{\beta,\kappa\}$. In the intermediate-frequency regime, since the RPL model typically overestimates the signal relative to the RPLP model, the constraints obtained with the RPL model would thus be conservative. In the context of environmental effects, equipped with a mapping of the parameters and a population, one can place upper bounds on specific environmental parameters. However, the RPL model can also be used to test several modified gravity theories with an appropriate mapping. This is similar to [47], where the authors used the parameterized post-Einsteinian (ppE) framework to place constraints on several modified theories of gravity. The advantage of our models over the ppE model of [47] is that we do not assume that the additional energy flux is small relative to the Newtonian GW flux. When the additional energy dissipation is small, results from our model will agree with those of [47]. To demonstrate our approach, we consider the case of a time-dependent Newton gravitational constant $G(t)$ (see [51, 90]) and obtain an order of magnitude upper bound on how the SGWB from sBBHs can place independent constraints on $|\dot{G}/G|$.
@@
A deviation from the vacuum GR prediction for the SGWB is constrainable if the turning point of the spectrum occurs in the detector sensitivity band. In Sec. IV.2.1, we found that using an astrophysical upper bound on the Eddington ratio with $f_{\rm Edd}\leq 100$, we could not constrain the effect of accretion, due to the turning point being at frequencies lower than the LISA sensitivity band. We now estimate the upper bound on $|\dot{G}_{0}/G_{0}|$ by requiring that the turning point of the signal be at $f\sim 10^{-3}$z, which is where LISA is optimally sensitive. We find that when $|\dot{G}_{0}/G_{0}|>10^{-4}\mathrm{yr}^{-1}$, the turning point would be above $10^{-3}$z, resulting in a $\gtrsim 10\%$ change in the vacuum GR signal, which can be tested with LISA. This order of magnitude constraint of $|\dot{G}_{0}/G_{0}|\lesssim 10^{-4}\mathrm{yr}^{-1}$ is nearly 10 orders of magnitude weaker than that from lunar ranging measurements [91] (see [92] for a review). However, as an independent constraint, it is comparable to what can be obtained by the loudest (and best known) verification Galactic binary in the LISA band [52], and slightly better than what neutron star mergers can constrain with future third-generation detectors [68]. Our estimate is also consistent with [49], when appropriately rescaling their results to mHz frequencies and stellar mass binaries. Thus, not only can the detection of SGWB of sBBHs in the LISA band probes environmental effects, but it also allows for independent and agnostic tests of GR, both of which can be accomplished using the phenomenological models developed in this work.
@@ VI Conclusion
The SGWB from sBBHs, which can be detected by LISA, offers a probe of the astrophysical environment in which these binaries form. This is enabled by a suppression of the SGWB due to additional energy-loss channels induced by the surrounding environment. In this work, we investigated the detectability (with LISA) of dynamical friction and gas accretion on the SGWB of sBBHs. Assuming that all sBBHs undergo either dynamical friction, our major results are that (i) for typical disk densities $\rho\sim 10^{-10}-10^{-9}\mathrm{g\,cm^{-3}}$, dynamical friction effects are well measured and the vacuum model is confidently disfavored even when including the Galactic foreground in the inference, (ii) gas accretion cannot inferred or constrained even for an Eddington ratio of $f_{\rm Edd}=10$ even excluding the Galactic foreground from the inference. As a consequence of (i), we also found that there are significant systematic biases in the inference of the vacuum amplitude and spectral index, highlighting the need to model environmental effects in the SGWB.
Further, we investigated the impact of a sBBH sub-population undergoing dynamical friction effects. As a proof-of-concept, using a mixture model, we showed that LISA is sensitive to dynamical friction effects (assuming a typical disk density) arising from a sBBH sub-population that contributes $10\%$ to the total SGWB. We constructed phenomenological parametric models that can capture environmental effects, and more generally, any additional energy loss channel. Thus our models can also be used to test GR, provided the modifications to GR are dominated by dissipative effects. As a proof-of-concept, we obtained an order-of-magnitude constraint on the time variation of Newton’s constant: $|\dot{G}/G|\lesssim 10^{-4}\mathrm{yr}^{-1}$, competitive with the projected constraint from the loudest LISA verification binary [52].
@@ Appendix A Merger rate and population model
[…]
@@
Recent observations indicate that the sBBH population exhibits low effective spins [61, 62]. Consequently, when computing the SGWB, we assume that sBBHs have negligible spin and focus solely on the mass distribution, which is described by the Power Law + Peak mode [59]. Under this assumption, the source parameters $\bm{\phi}$ in Eq. 2 contain the primary mass $m_{1}$ and mass ratio $q$, and the distribution $p(\bm{\phi})$ reduces to joint probability density function $p(m_{1},q|~{}{\mathrm{\Lambda}_{m}})$, where ${\mathrm{\Lambda}_{m}}=\left\{\alpha_{\rm pop},\beta_{\rm pop},m_{\min},m_{\max},\delta_{\rm m},\lambda_{\rm peak},\mu_{\rm m},\sigma_{\rm m}\right\}$ represents the set of hyperparameters governing the mass distribution.
@@
We numerically evaluate Eq. 2 via Monte Carlo integration, using the corresponding posterior distributions of merger rate and mass distribution models inferred from [43, 5] and generate posterior predictions for the SGWB. We obtain the predictive posterior median by evaluating, at each frequency, the median of the SGWB realizations from the posterior samples. By computing the $\ell^{2}$-norm between each sampled SGWB spectrum and the median curve, we select the sample that minimizes this distance and the corresponding parameters are adopted as fiducial values in this work. The merger rate parameters are taken as $R_{0}=21~{}\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, $\lambda_{1}=1.5$, $\lambda_{2}=3.7$, and $z_{p}=3.0$. The hyperparameters of mass distribution are $\alpha_{\rm pop}=3.6$, $\beta_{\rm pop}=3.4$, $m_{\min}=4.6M_{\odot}$, $m_{\max}=84M_{\odot}$, $\delta_{\rm m}=4.5M_{\odot}$, $\lambda_{\rm peak}=0.018$, $\mu_{\rm m}=29,M_{\odot}$, and $\sigma_{\rm m}=8.3,M_{\odot}$.
In Fig. 10, we demonstrate the consistency between the median predicted SGWB curve generated by our adopted model and the forecast based on the O3 observational data [5]. The remapped spectrum corresponding to the fiducial parameter set closely tracks the predictive median, and both lie well within the $90\%$ credible interval of the O3 inference. Further, we observe that the spectral indices match well at low frequencies corresponding to the inspiral regime. However, at higher frequencies, higher order PN effects become relevant as the spectrum transitions toward merger and ringdown regimes. Since we use a leading PN approximation to compute the energy sepctrum given by Eq. 3, we find that the spectral index is overestimated relative to the O3 inference that uses the higher PN terms. In the lower-frequency bands relevant to LISA, which is our primary focus, the energy spectrum of sBBHs is well described by the leading PN expression of Eq. 3, and thus we are justified in neglecting higher PN terms for this work.
@@ B.1 Validity of dynamical friction modeling
Requiring the dynamical friction force to be perturbative is equivalent to requiring that the dynamical friction timescale $\tau_{\rm DF}$ is much longer than the Keplerian orbital timescale $\tau_{\rm orb}$. In a multiple-scale-analysis treatment [69], we essentially consider dynamical friction to leading order in $\epsilon_{\rm DF}\sim\tau_{\rm orb}/\tau_{\rm DF}$. We can see that $\epsilon_{\rm DF}$ scales with $v^{-6}$, making it a $-3$PN relative contribution to the Keplerian equations of motion. However, being a dissipative effect, the dynamical friction force scales as $v^{-11}$ relative to the GW radiation reaction force, making it a $-5.5$PN contribution at the level of the GW flux. The important point here is that the perturbative $\epsilon_{\rm DF}\ll 1$ regime does not necessarily imply that dynamical friction is perturbative relative to GW radiation reaction, as is typically assumed when computing the GW phase [25].
For a typical stellar binary with comparable masses, in the very early inspiral regime of $f_{s}\sim 10^{-5}$z, dynamical friction will become non-perturbative when $\epsilon_{\rm DF}\sim 1$, resulting in
@@
We find that the agnostic upper bound from Eq. 34 is consistent with typical densities at the migration trap of thin Keplerian accretion disks around supermassive black holes (SMBHs) as discussed in [19]. Thus, for densities with $\rho\ll\rho_{\rm upper,modeling}$, using linear perturbation theory is sufficient. In our analysis, we consider values of $\rho\lesssim 10^{-7}\mathrm{g}\,\mathrm{cm}^{-3}$, which corresponds to $\epsilon_{\rm DF}\lesssim 10^{-2}$ that shows that we are well within the perturbative regime.
@@
Furthermore, relativistic effects such as 1PN corrections to the dynamical friction evolution (formally scaling as $\mathcal{O}(\rho v^{2})$ relative to the Keplerian dynamics) will also become increasingly relevant as the binary inspirals. These relativistic corrections are relevant for predicting the long-time evolution of a single binary, e.g. the evolution of eccentricity as shown in the toy model of [18]. However, for predicting the SGWB, at the frequencies where Eq. 6 breaks down or mutual wake effects or relativistic couplings become relevant, GW emission will completely dominate the effect of dynamical friction, simply due to the relative frequency scaling of $f^{-11/3}$. Essentially, at higher frequencies the GW energy spectrum will rapidly asymptote to the vacuum SGWB, and higher order modeling of dynamical friction effects (that only matter at high frequencies) will not affect the prediction of the SGWB. The main impact of dynamical friction is at lower frequencies, where Eq. 6 is valid, and mutual wake effects and relativistic couplings can be safely neglected.

### v2

@@ IV.3 Measurability of dynamical friction effects from a sub-population of sBBHs
Figure 8: SGWB spectra $\Omega_{\mathrm{GW}}(f)$ for various environmental fractions $p_{\mathrm{env}}$ in the case of dynamical friction with $\rho^{\rm inj}=10^{-10}\mathrm{g\,cm^{-3}}$. The gray dashed and black dotted curves indicate the power-law sensitivity (PLS) and Bayesian power-law sensitivity (BPLS) of LISA, under the same assumptions in Figs.˜2 and 1. Here, BF denotes the Bayes factor of BPLS.
The impact of the SGWB due to environmental effects will depend on the distribution of the environmental parameters ($\rho$ and $f_{\rm Edd}$) across the population. In a realistic scenario, only a sub-population of sBBHs formed in gaseous medium, where the environmental effects are significant. As a proof-of-concept, we phenomenologically construct a mixture model to account for environmental effects arising from a sub-population. We introduce the fraction parameter $p_{\mathrm{env}}$, which characterizes the fractional contribution of the SGWB arising from binaries affected by environmental effects. The total SGWB is then given by
@@
In Fig.˜8, for a given injected density of $\rho^{\rm inj}=10^{-10}\mathrm{g}\,\mathrm{cm}^{-3}$, we illustrate how changing the injected fraction $p_{\rm env}^{\rm inj}$ leads to changes in the SGWB. The case of $p_{\rm env}=0$ is identical to the vacuum SGWB, while $p_{\rm env}=1$ is identical to having all binaries affected by environmental effects (as shown in Fig.˜1). Notably, for $0<p_{\rm env}<1$ the total SGWB exhibits a spectral shape resulting from the $p_{\rm env}$-weighted superposition of two SGWBs, dominated by vacum component at low frequencies. If future LVK observations provide tighter constraints on the binary population, the difference in amplitude of the SGWB at high and low frequencies might signal the presence of a sub-population containing environmental effects.
@@
With only a sub-population affected by environmental effects, we assess how the measurability of $\rho$ changes with $p_{\rm env}$. To do so, we first generate different injections of $\Omega_{\rm Frac}$ using Eq.˜29 for $p^{\rm inj}_{\rm env}\in\{0.1,0.3,0.5,0.7,0.9\}$ and a fixed density $\rho^{\rm inj}=10^{-10}\mathrm{g}\,\mathrm{cm}^{-3}$. Specifically, in Eq.˜29, we evaluate $\Omega_{\rm env}$ and $\Omega_{\rm vac}$ as integrals over the population, which we discussed in Sec. III.1 and Sec. II respectively. We then perform parameter estimation using $\Omega_{\rm Frac}$, but with $\Omega_{\rm env}$ and $\Omega_{\rm vac}$ evaluated using the RPLP model. Note that the RPLP model describes $\Omega_{\rm vac}$ with $\alpha=0$. As in Sec. IV.2.1, we perform the analysis with and without the Galactic foreground.
In Fig.˜9, the blue (orange) histograms correspond to the one-dimensional marginalized posteriors of $\rho$ with (without) including the Galactic foreground parameters. The first five panels of Fig.˜9 correspond to the cases with different injected fraction values. Meanwhile, for reference, the last panel of Fig.˜9 shows the case where the injected data contains dynamical friction effects for all binaries with $p_{\rm env}=1$, which we had shown earlier in Fig.˜3. We find that (for this injected density) the impact of the injected fraction on the marginalized posterior of $\rho$ is not significant when including the Galactic foreground. This is due to the fact that the Galactic foreground is not suppressed in the regime where the signal undergoes a turning point as a result of dynamical friction. When excluding the Galactic foreground, we can better see the impact of the sub-population on the marginalized posterior of $\rho$.
For each injection, we also separately infer with the vacuum model and compute the log Bayes factor $\log_{10}\mathcal{B}^{\rm frac}_{\rm vac}$ between the mixture and vacuum models. We find that for all values of the injected fraction, there is substantial preference for the mixture model over the vacuum model with Bayes factors in the range $\log_{10}\mathcal{B}^{\rm Frac}_{\rm vac}\sim 0.79$$0.95$ depending on the fraction. When excluding the Galactic foreground, we instead have $\log_{10}\mathcal{B}^{\rm Frac}_{\rm vac}\sim 1.27$$3.52$ depending on the fraction, indicating strong preference for the mixture model. In other words, the evidence in favor of the mixture model (over vacuum) is weaker when including the Galactic foreground, as we had also seen in previous analysis with $p_{\rm env}=1$. We have thus shown that even with only sub-populations undergoing dynamical friction effects, and with including the Galactic foreground, we can reject the vacuum model with substantial evidence.
@@
Figure 9: Marginalized posterior distribution of $\log_{10}(\rho/\rho_{0})$ as a function of the environmental fraction parameter $p_{\rm env}$, with an injected density $\rho^{\rm inj}=10^{-10}\rm{g\,cm^{-3}}$ (horizontal dashed line). The first five panels show the recovered posteriors using the mixture model for different injected values of $p_{\rm env}$. The last panel is for reference and shows the recovered posteriors using the RPLP model when injecting with $p_{\rm env}=1$ (as shown in Fig.˜3).
@@ V Agnostic tests of additional energy loss channels
In general, when there is energy loss from binaries in addition to GW emission, the SGWB will decrease below the vacuum GR prediction given by Eq.˜3. We use our phenomenological model to probe such additional energy losses agnostically. Specifically, with our RPL model (see Eq.˜17), one can constrain the dimensional combination $\tilde{\alpha}\equiv A_{m}\alpha$ for different choices of $\{\beta,\kappa\}$. In the intermediate-frequency regime, since the RPL model typically overestimates the signal relative to the RPLP model, the constraints obtained with the RPL model would thus be conservative. In the context of environmental effects, equipped with a mapping of the parameters and a population, one can place upper bounds on specific environmental parameters. However, the RPL model can also be used to test several modified gravity theories with an appropriate mapping. This is similar to [55], where the authors used the parameterized post-Einsteinian (ppE) framework to place constraints on several modified theories of gravity. The advantage of our models over the ppE model of [55] is that we do not assume that the additional energy flux is small relative to the Newtonian GW flux. When the additional energy dissipation is small, results from our model will agree with those of [55]. To demonstrate our approach, we consider the case of a time-dependent Newton gravitational constant $G(t)$ (see [59, 98]) and obtain an order of magnitude upper bound on how the SGWB from sBBHs can place independent constraints on $|\dot{G}/G|$.
@@
A deviation from the vacuum GR prediction for the SGWB is constrainable if the turning point of the spectrum occurs in the detector sensitivity band. In Sec. IV.2.1, we found that using an astrophysical upper bound on the Eddington ratio with $f_{\rm Edd}\leq 100$, we could not constrain the effect of accretion, due to the turning point being at frequencies lower than the LISA sensitivity band. We now estimate the upper bound on $|\dot{G}_{0}/G_{0}|$ by requiring that the turning point of the signal be at $f\sim 10^{-3}$Hz, which is where LISA is optimally sensitive. We find that when $|\dot{G}_{0}/G_{0}|>10^{-4}\mathrm{yr}^{-1}$, the turning point would be above $10^{-3}$Hz, resulting in a $\gtrsim 10\%$ change in the vacuum GR signal, which can be tested with LISA. This order of magnitude constraint of $|\dot{G}_{0}/G_{0}|\lesssim 10^{-4}\mathrm{yr}^{-1}$ is nearly 10 orders of magnitude weaker than that from lunar ranging measurements [99] (see [100] for a review). However, as an independent constraint, it is comparable to what can be obtained by the loudest (and best known) verification Galactic binary in the LISA band [60], and slightly better than what neutron star mergers can constrain with future third-generation detectors [75]. Our estimate is also consistent with [57], when appropriately rescaling their results to mHz frequencies and stellar mass binaries. Thus, not only can the detection of SGWB of sBBHs in the LISA band probes environmental effects, but it also allows for independent and agnostic tests of GR, both of which can be accomplished using the phenomenological models developed in this work.
@@ VI Conclusion
The SGWB from sBBHs, which can be detected by LISA, offers a probe of the astrophysical environment in which these binaries form. This is enabled by a suppression of the SGWB due to additional energy-loss channels induced by the surrounding environment. In this work, we investigated the detectability (with LISA) of dynamical friction and gas accretion on the SGWB of sBBHs. Assuming that all sBBHs undergo either dynamical friction or gas accretion, our major results are that (i) for typical disk densities $\rho\sim 10^{-10}-10^{-9}\mathrm{g\,cm^{-3}}$, dynamical friction effects are well measured and the vacuum model is confidently disfavored even when including the Galactic foreground in the inference, (ii) gas accretion cannot inferred or constrained even for an Eddington ratio of $f_{\rm Edd}=10$ even excluding the Galactic foreground from the inference. As a consequence of (i), we also found that there are significant systematic biases in the inference of the vacuum amplitude and spectral index, highlighting the need to model environmental effects in the SGWB.
Further, we investigated the impact of a sBBH sub-population undergoing dynamical friction effects. As a proof-of-concept, using a mixture model, we showed that even when only a sub-population of sBBHs undergoes dynamical friction with moderate values of the disk density, LISA can substantially reject the vacuum model, while still including the Galactic foreground parameters. We constructed phenomenological parametric models that can capture environmental effects, and more generally, any additional energy loss channel. Thus our models can also be used to test GR, provided the modifications to GR are dominated by dissipative effects. As a proof-of-concept, we obtained an order-of-magnitude constraint on the time variation of Newton’s constant: $|\dot{G}/G|\lesssim 10^{-4}\mathrm{yr}^{-1}$, competitive with the projected constraint from the loudest LISA verification binary [60].
@@ Appendix A Merger rate and population model
$\tilde{R}_{\mathrm{GW}}(z)=\int dt_{d}R_{\mathrm{SFR}}\left(z_{f}\right)F\left(Z<Z_{\mathrm{thresh}},z_{f}\right)p\left(t_{d}\right),$ (31)
@@
Recent observations indicate that the sBBH population exhibits low effective spins [66, 67]. Consequently, when computing the SGWB, we assume that sBBHs have negligible spin and focus solely on the mass distribution, which is described by the Power Law + Peak mode [64]. Under this assumption, the source parameters $\bm{\phi}$ in Eq.˜2 contain the primary mass $m_{1}$ and mass ratio $q$, and the distribution $p(\bm{\phi})$ reduces to joint probability density function $p(m_{1},q|~{\mathrm{\Lambda}_{m}})$, where ${\mathrm{\Lambda}_{m}}=\left\{\alpha_{\rm pop},\beta_{\rm pop},m_{\min},m_{\max},\delta_{\rm m},\lambda_{\rm peak},\mu_{\rm m},\sigma_{\rm m}\right\}$ represents the set of hyperparameters governing the mass distribution.
@@
We numerically evaluate Eq.˜2 via Monte Carlo integration, using the corresponding posterior distributions of merger rate and mass distribution models inferred from [51, 5] and generate posterior predictions for the SGWB. We obtain the predictive posterior median by evaluating, at each frequency, the median of the SGWB realizations from the posterior samples. By computing the $\ell^{2}$-norm between each sampled SGWB spectrum and the median curve, we select the sample that minimizes this distance and the corresponding parameters are adopted as fiducial values in this work. The merger rate parameters are taken as $R_{0}=21~\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, $\lambda_{1}=1.5$, $\lambda_{2}=3.7$, and $z_{p}=3.0$. The hyperparameters of mass distribution are $\alpha_{\rm pop}=3.6$, $\beta_{\rm pop}=3.4$, $_{\min}=4.6M_{\odot}, mm$$_{\max}=84M_{\odot}$, $\delta_{\rm m}=4.5M_{\odot}$, $\lambda_{\rm peak}=0.018$, $\mu_{\rm m}=29,M_{\odot}$, and $\sigma_{\rm m}=8.3,M_{\odot}$.
In Fig.˜10, we demonstrate the consistency between the median predicted SGWB curve generated by our adopted model and the forecast based on the O3 observational data [5]. The remapped spectrum corresponding to the fiducial parameter set closely tracks the predictive median, and both lie well within the $90\%$ credible interval of the O3 inference. Further, we observe that the spectral indices match well at low frequencies corresponding to the inspiral regime. However, at higher frequencies, higher order PN effects become relevant as the spectrum transitions toward merger and ringdown regimes. Since we use a leading PN approximation to compute the energy sepctrum given by Eq.˜3, we find that the spectral index is overestimated relative to the O3 inference that uses the higher PN terms. In the lower-frequency bands relevant to LISA, which is our primary focus, the energy spectrum of sBBHs is well described by the leading PN expression of Eq.˜3, and thus we are justified in neglecting higher PN terms for this work.
@@ B.1 Validity of dynamical friction modeling
Requiring the dynamical friction force to be perturbative is equivalent to requiring that the dynamical friction timescale $\tau_{\rm DF}$ is much longer than the Keplerian orbital timescale $\tau_{\rm orb}$. In a multiple-scale-analysis treatment [76], we essentially consider dynamical friction to leading order in $\epsilon_{\rm DF}\sim\tau_{\rm orb}/\tau_{\rm DF}$. We can see that $\epsilon_{\rm DF}$ scales with $v^{-6}$, making it a 3-3PN relative contribution to the Keplerian equations of motion. However, being a dissipative effect, the dynamical friction force scales as $v^{-11}$ relative to the GW radiation reaction force, making it a 5.5-5.5PN contribution at the level of the GW flux. The important point here is that the perturbative $\epsilon_{\rm DF}\ll 1$ regime does not necessarily imply that dynamical friction is perturbative relative to GW radiation reaction, as is typically assumed when computing the GW phase [26].
For a typical stellar binary with comparable masses, in the very early inspiral regime of $f_{s}\sim 10^{-5}$Hz, dynamical friction will become non-perturbative when $\epsilon_{\rm DF}\sim 1$, resulting in
@@
We find that the agnostic upper bound from Eq.˜34 is consistent with typical densities at the migration trap of thin Keplerian accretion disks around supermassive black holes (SMBHs) as discussed in [20]. Thus, for densities with $\rho\ll\rho_{\rm upper,modeling}$, using linear perturbation theory is sufficient. In our analysis, we consider values of $\rho\lesssim 10^{-7}\mathrm{g}\,\mathrm{cm}^{-3}$, which corresponds to $\epsilon_{\rm DF}\lesssim 10^{-2}$ that shows that we are well within the perturbative regime.
@@
Furthermore, relativistic effects such as 1PN corrections to the dynamical friction evolution (formally scaling as $\mathcal{O}(v^{2}\epsilon_{\rm DF})$ relative to the Keplerian dynamics) will also become increasingly relevant as the binary inspirals. These relativistic corrections are relevant for predicting the long-time evolution of a single binary, e.g. the evolution of eccentricity as shown in the toy model of [18]. However, for predicting the SGWB, at the frequencies where Eq.˜6 breaks down or mutual wake effects or relativistic couplings become relevant, GW emission will completely dominate the effect of dynamical friction, simply due to the relative frequency scaling of $f^{-11/3}$. Essentially, at higher frequencies the GW energy spectrum will rapidly asymptote to the vacuum SGWB, and higher order modeling of dynamical friction effects (that only matter at high frequencies) will not affect the prediction of the SGWB. The main impact of dynamical friction is at lower frequencies, where Eq.˜6 is valid, and mutual wake effects and relativistic couplings can be safely neglected.